OpenStudy (anonymous):

solve for t (5/13)t = -9

jimthompson5910 (jim_thompson5910):

(5/13)t = -9 5t = -9*13 5t = -117 t = -117/5

OpenStudy (anonymous):

thank you very much sir thompson